Consider the following information developed by the accountant at ABC company that produces and sells leather cases: Total Sales (10.000 units) LE 4.000.000 COGS 2.400.000 Selling and administrative expenses 1.500.000 Other information: 1- Manufacturing variable cost per unit LE 160. 2- Selling and administrative fixed cost LE 700.000 Requires: 1- Present the contribution margin income statement (8 marks). 2- Calculate break-even point in units (3 marks) 3- Calculate break-even point value (3 marks) 4- Calculate budgeted sales in units that should be reached to earn profit of LE 300.000 (3 marks) 5- Calculate budgeted sales value that should be reached to earn profit of LE 400.000 (3 marks) 6- Suppose the target net income after taxes was LE 500.000 at income tax rate of 20%, calculate the following: a- Target net income before taxes. (2 marks) b- Target sales in units. (4 marks) c- Suppose target net income after taxes was LE 625.000, calculate the targeted sales value. (4 marks


1
Expert's answer
2021-05-31T15:50:08-0400

Solution:

1.). Contribution margin income statement

See below:




 

2.). Break-even point in units:

Break-even point in units = Fixed Costs ÷ (Revenue per Unit - Variable Cost per Unit)

Break-even point in units = "\frac{700,000}{(400 - 160)} = 2,917 units"


3.). Break-even in value:

Break-even in value = "\frac{Fixed \;Costs}{Contribution \;margin}"


Contribution margin = "\frac{Contribution}{Total \;sales}" "=\frac{100,000}{4,000,000}\times 100 = 2.5\%"


Break-even in value = "=\frac{700,000}{2.5\%} = 28,000,000"


4.). Total sales – COGS – Selling and administrative expenses – Fixed costs = 300,000

X – 2,400,000 – 1,500,000 – 700,000 = 300,000

X – 4,600,000 = 300,000

X = 300,000 + 4,600,000 = 4,900,000

Total Sales = 4,900,000

Budget sales in units = 4,900,000/10,000 = 490 units

5.). Total sales – COGS – Selling and administrative expenses – Fixed costs = 400,000

X – 2,400,000 – 1,500,000 – 700,000 = 400,000

X – 4,600,000 = 400,000

X = 300,000 + 4,600,000 = 5,000,000

Total Sales = 5,000,000


6.). a). Target net income before taxes:

Net income after taxes is 80%, find 100%?

Net income before taxes = "500,000\times \frac{100\%}{80\%} = 625,000"

Net income before taxes = LE 625,000


b.). Target sales in units = "\frac{625,000}{10,000} = 62.5 \;units"


c.). Targeted sales value:

First derive net income before taxes:

Net income after taxes is 80%, find 100%?

Net income before taxes = "625,000\times \frac{100\%}{80\%} = 781,250" 6

Total sales – COGS – Selling and administrative expenses – Fixed costs = 781,250

X – 2,400,000 – 1,500,000 – 700,000 = 781,250

X – 4,600,000 = 781,250

X = 781,250 + 4,600,000 = 1,241,250

Total Target Sales = 1,241,250